Quad Enterprises is considering a new three-year expansion project that requires an initial fixed asset investment of $2.31 million. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $1,657,000 in annual sales, with costs of $633,000. If the tax rate is 25 percent, what is the OCF for this project? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, e.g., 1,234,567.)
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(2,310,000/3)=$770,000/year

Hence OCF=(Sales-Costs)(1-tax rate)+Tax savings on Annual depreciation

=(1,657,000-633,000)(1-0.25)+(0.25*770,000)

which is equal to

=$960,500